OpenStudy (anonymous):

d/dx(arctan((x+3)(2x-2)))=?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

chain rule for this one

OpenStudy (misty1212):

\[\frac{1}{((x+3)(2x-2))^2}\times \frac{d}{dx}[(x+3)(2x-2)]\]

OpenStudy (misty1212):

ok that was wrong

OpenStudy (misty1212):

\[\frac{1}{((x+3)(2x-2))^2+1}\times \frac{d}{dx}[(x+3)(2x-2)]\]

OpenStudy (anonymous):

thank you